The electrical circuit has 3 devices each plugged into household powerpoint. Device A has a resistance of 1.0 Ω, device B has a resistance of 3.0 Ω and device C has a resistance of 6.0 Ω.

If all three devices are turned on for 15 minutes, what will be the Total Energy Consumption of the 3 device.
1
Expert's answer
2020-05-04T12:41:13-0400

if we look at the connections of the resistances then it is a parallel combination

hence the equivalent resistance is "1||3||6" =2/3 ohm

suppose the potential provide was V volts

then the power = "V^2\/R=2V^2\/3"

energy consumption = "(2V^2/3)\ \times15\times60=600V^2"